I am estimating how much I would need to charge for painting 148 rooms in a new motel each room measuring 14 ft by 16 ft by 10 ft high. What is the area?
1
Expert's answer
2013-03-13T06:48:57-0400
Perimeter of a room=(length + width)*2=(14+16)*2=60(ft)
Area of walls& of& a single room=perimeter*height=60*10=600(sq.ft ).
Area& of ceiling& of a single room=length*width=14*16=224 (sq.ft).
Area for painting in a single room= Area of walls + Area of a ceiling=600+224=824(sq.ft).
Total area in a motel for painting 148 rooms = 824*148=121952 (sq.ft).
Answer: 121952 (sq.ft).